# From Word Vectors to Multimodal Embeddings: Techniques, Applications, and Future Directions For Large Language Models

### v1

@@ II-E1 Modeling Individual Linguistic Variation
Personalized word embeddings aim to capture individual differences in word usage and linguistic preferences, advancing beyond standard generic representations. These embeddings can enhance language model performance in tasks customized to specific users. [31] investigated the value of personalized word embeddings in language modeling, finding that combining generic and personalized embeddings led to a 4.7% relative reduction in perplexity, thus improving model performance. They also observed that words associated with specific psycholinguistic categories showed greater variation across users, indicating that personalized models are particularly advantageous for predicting such words. Additionally, [32] examined properties encoded within sentence embeddings, providing insights into the influence of factors such as word frequency and positional distance on a model’s encoding of content and word order. Their findings on individual linguistic variation at the sentence level offer valuable guidance for developing more user-specific word embeddings and NLP systems.
@@ II-E2 Applications of Personalized Embeddings
Personalized word embeddings have proven beneficial across various NLP tasks. [31] demonstrated the potential of personalized embeddings in authorship attribution, showing that they can effectively capture unique writing styles. [11] introduced LSTM-RNN models for sentence embedding trained on user click-through data, underscoring the application of personalized embeddings in user-specific document retrieval systems that provide more relevant search results and tailored recommendations. Personalized embeddings can also leverage user history data: [33] proposed a User Embedding Module (UEM) that compresses user histories into embeddings to serve as prompts in language models. This approach yielded improvements in personalized language tasks, suggesting that personalized embeddings effectively capture user preferences and interests for customized recommendations and user profiling.
@@ III-A1 Simple Averaging and Pooling of Word Embeddings
  • Methods: Common methods include averaging the word embeddings, or using element-wise maximum (max-pooling), minimum (min-pooling), or sum operations. These approaches have been explored in various contexts, such as in a comparative study of word embedding models for hate speech detection [47].
  • Strengths: The main advantage of these methods is their simplicity and computational efficiency. They are easy to implement and require minimal computational resources, making them ideal for large-scale applications. For example, a study comparing bag-of-words approaches with sentence encoders found that simple averaging can be unexpectedly effective for certain tasks [48].
  • Limitations: However, these methods have notable limitations. Averaging or pooling word embeddings typically results in a loss of word order information, which limits their effectiveness in tasks where word order is significant. Additionally, they can be sensitive to outliers, as an uncommon word may disproportionately influence the resulting sentence embedding. To mitigate these issues, researchers have proposed more advanced techniques, such as using linear combinations of word embeddings to capture higher-level linguistic structures [49].
@@ III-A2 Recurrent Neural Network (RNN) based Approaches
  • Architectures: LSTM-RNNs and GRUs are commonly used for sentence embedding as they effectively capture long-range dependencies within sentences. These architectures surpass simple averaging methods in tasks that require deeper contextual understanding. For instance, a study on deep sentence embedding with LSTM networks demonstrated their capacity to detect keywords and allocate topics autonomously [11].
  • Training: RNN-based models can be trained using various approaches, including weakly supervised methods. One example involves training LSTM-RNNs with user click-through data from web search engines, which helps models learn effective representations for document retrieval [11].
  • Applications: RNN-generated sentence embeddings have been applied to numerous NLP tasks, such as document retrieval and semantic similarity measurement. For instance, using LSTM-RNN embeddings for document retrieval yielded notable performance improvements over traditional methods [11].
@@ III-A3 Transformer-based Sentence Encoders
Sentence-BERT (SBERT) has become a leading architecture for generating sentence embeddings, as highlighted by [50]. SBERT leverages pre-trained transformer models like BERT, fine-tuning them specifically for sentence-level tasks to generate efficient and effective sentence embeddings. Another notable model, SBERT-WK [51], refines BERT-based word models to produce high-quality sentence representations. This model focuses on analyzing word representation evolution across BERT layers, applying geometric analysis to enhance sentence embeddings.
@@
Fine-tuning strategies are essential for adapting pre-trained language models to sentence embedding tasks. A common method is supervised fine-tuning on Natural Language Inference (NLI) and Semantic Textual Similarity (STS) datasets, as discussed in [29]. This approach uses labeled sentence pairs with semantic relationship annotations (e.g., entailment, contradiction, similarity) to refine model performance. Knowledge distillation also proves effective; [52] demonstrates how a student model can be trained to mimic a teacher model, effectively transferring insights from a high-resource language to multiple target languages.
@@
Transformer-based sentence encoders achieve state-of-the-art results in various Semantic Textual Similarity (STS) tasks [53]. These models excel at capturing semantic relationships, making them highly effective for tasks requiring understanding of sentence meaning and similarity. Specifically, SBERT-WK demonstrates superior performance on STS benchmarks, further validating its robustness in sentence embedding applications [51].
@@
Analyzing the evolution of word representations across transformer layers provides insights into how these models capture and process contextual information. Studies such as [54] introduce methods to quantify and visualize contextualization across layers, often analyzing changes in word embeddings and examining the unique information captured by each layer. Additionally, [55] investigates word representations at both input and output stages, offering deeper insights into layer-wise contextual processing.
@@
The BeLLM model [56] addresses the limitations of unidirectional attention in autoregressive large language models by introducing backward dependencies. By transforming specific attention layers to be bidirectional, BeLLM enhances semantic similarity tasks by considering both preceding and following context, thereby providing a more comprehensive contextual understanding.
@@ III-A4 Multilingual and Cross-lingual Sentence Embeddings
Knowledge distillation is an effective technique for aligning vector spaces across languages, facilitating cross-lingual understanding [52]. By training a student model to replicate the vector space of a monolingual teacher model, semantic knowledge from a high-resource language can be transferred to other languages, including those with limited resources.
@@
Dual encoder architectures and negative sampling are core techniques in translation ranking and bi-text mining tasks, as explored by [29]. In these architectures, source and target sentences are encoded separately, with training focused on ranking true translation pairs higher than negative samples, thereby improving translation accuracy and relevance.
@@
The Language-agnostic BERT Sentence Embedding (LaBSE) model [29] has demonstrated high performance in multilingual applications. LaBSE utilizes a combination of pre-training and dual-encoder fine-tuning to optimize translation ranking and produce language-agnostic sentence embeddings that excel across languages.
@@
Addressing language bias is essential to ensure fair and effective multilingual embeddings. [52] notes that some models may exhibit biases, favoring specific language combinations over others, which can impair performance in multilingual tasks. Techniques such as knowledge distillation are valuable for mitigating this bias and achieving more balanced cross-lingual representations.
@@
Handling low-resource languages remains challenging in multilingual NLP. [28] highlights that bilingual lexicons can provide an effective solution by enabling cross-lingual word embedding generation when text data is scarce. This approach enhances language model performance in low-resource environments, helping bridge the gap in multilingual understanding.
@@ III-B1 Methods for Combining Word Embeddings
A straightforward approach to creating document embeddings is to average the word embeddings of all words in the document. Although simple, this method can be highly effective and sometimes outperforms more complex models, particularly when word order is not critical [49]. Averaging captures semantic relationships by leveraging word co-occurrence patterns [49]. However, this method overlooks word order and may miss nuanced information in complex sentences [49].
Pooling methods, such as max-pooling or min-pooling, offer an alternative by selecting the most salient features from word embeddings. While these methods can highlight important features, they may still lose significant contextual information. More sophisticated techniques, like using recurrent neural networks (RNNs) with Long Short-Term Memory (LSTM) cells, process words sequentially, accumulating richer information as they traverse the document [11]. This method has shown effectiveness in tasks like document retrieval, as demonstrated in [11].
@@
Beyond simple averaging and pooling, compositional methods aim to capture interactions between words within a document. [57] explores compositional embedding models that create representations for substructures within annotated sentences, addressing the challenge of representing a diverse range of sentences by building sentence embeddings from component word embeddings. Additionally, [7] presents an approach that learns embeddings for phrases by combining word embeddings, providing richer phrase-level representations for documents.
@@ III-B2 Generative Topic Embedding Models
Integrating word embeddings into topic models, such as Latent Dirichlet Allocation (LDA), enhances the quality of learned topics. [44] proposes a generative topic embedding model that merges local word collocation patterns with global document-level topic patterns through variational inference. This approach yields topic embeddings and topic mixing proportions for each document, resulting in low-dimensional, continuous representations [44].
@@
Correlated topic models aim to capture relationships between topics. [58] introduces the Correlated Gaussian Topic Model (CGTM), which uses word embeddings to represent topics as multivariate Gaussian distributions. This structure captures semantic relatedness and correlations between words, improving topic coherence over traditional topic models [58].
@@
The Embedded Topic Model (ETM), described in [59], combines word embeddings with traditional topic models to improve topic quality and predictive performance. Unlike LDA, ETM uses word and topic embeddings to model word probabilities within a topic, effectively handling large vocabularies and rare words. This model has demonstrated superior performance over LDA in settings with extensive vocabularies [59].
@@ III-B3 Hierarchical Neural Language Models for Documents
Hierarchical neural models often use layered structures to capture different aspects of context. [12] presents a two-layer model: the upper layer models the temporal context of document sequences, assuming that temporally close documents are likely to be related, while the lower layer captures word sequence context within each document. This setup enables the model to learn joint representations for both documents and words within a shared feature space, which simplifies tasks like document tagging through nearest-neighbor search. This hierarchical structure can be extended with additional layers, such as a user layer, to incorporate user preferences for personalized recommendations.
@@
A key advantage of hierarchical neural models is their ability to incorporate temporal context. [12] shows how these models learn representations that account for temporal co-occurrences of documents within a stream. By optimizing the joint log-likelihood of document and word sequences, the model balances context at both levels, moving beyond traditional bag-of-words approaches. This temporal awareness supports a more dynamic understanding of language, resulting in improved performance in tasks like movie classification (MovieLens dataset) and click-through prediction (Yahoo News dataset).
@@
The joint representations of documents and words learned by hierarchical neural language models enable a range of applications. For document retrieval, the similarity between a query and document can be efficiently computed by comparing their vector representations in the shared embedding space. In recommendation systems, the model leverages user preferences, represented in an additional user layer, to generate personalized suggestions. For automatic tagging, the shared embedding space facilitates nearest-neighbor searches to assign relevant tags based on document vectors. [12] demonstrates the model’s superior performance on tasks like movie classification, underscoring its versatility in document retrieval and recommendation. The ability to capture both document- and word-level context makes these models highly applicable for text mining.
{forest}
[…]
and
Document
Embeddings, root, [Cross-Lingual and
Multi-Lingual
Embeddings, level1, [Cross-Lingual Word Embeddings [30, 28], level2] [Challenges and Evaluation [60, 61], level2] [Language-Agnostic Sentence Embeddings [52, 29], level2] ] [Document Embeddings, level1, [Methods for Combining Word Embeddings
[49, 11, 57, 7], level2] [Generative Topic Embedding Models [44, 58, 59], level2] [Hierarchical Neural Language Models [12], level2] ] [Sentence Embeddings, level1, [Simple Averaging and Pooling [47, 48, 49], level2] [Recurrent Neural Networks [11], level2] [Transformer-based Sentence Encoders
[50, 51, 29, 52, 54, 55, 56], level2] ] ]
@@ III-C1 Cross-Lingual Word Embeddings: Methods and Applications
Cross-lingual word embeddings map words from different languages into a shared vector space, facilitating cross-lingual knowledge transfer. This approach is especially useful for low-resource languages, which often lack sufficient monolingual training data. Supervised methods typically use bilingual lexicons or word alignments from parallel corpora. For example, [30] introduces bilingual word embeddings learned from a large, unlabeled corpus, where machine translation word alignments are used to enforce translational equivalence. These embeddings, when integrated into a phrase-based machine translation system, show significant improvements in translation quality. Another supervised approach, explored in [28], leverages bilingual lexicons to enhance language models for low-resource languages, applying this technique to the language Yongning Na.
In contrast, unsupervised methods develop cross-lingual embeddings without explicit parallel data. [60] proposes a multilingual neural language model that shares parameters across languages, jointly learning word embeddings in a shared space without parallel data or pre-training. This method proves effective even with limited monolingual data or domain differences across languages. For a comprehensive overview of contextual and cross-lingual embeddings, [62] offers valuable insights.

### v2

@@ II-E1 Modeling Individual Linguistic Variation
Personalized word embeddings aim to capture individual differences in word usage and linguistic preferences, advancing beyond standard generic representations. These embeddings can enhance language model performance in tasks customized to specific users. [charles_exploring_2020] investigated the value of personalized word embeddings in language modeling, finding that combining generic and personalized embeddings led to a 4.7% relative reduction in perplexity, thus improving model performance. They also observed that words associated with specific psycholinguistic categories showed greater variation across users, indicating that personalized models are particularly advantageous for predicting such words. Additionally, [yossi_analysis_2017] examined properties encoded within sentence embeddings, providing insights into the influence of factors such as word frequency and positional distance on a model’s encoding of content and word order. Their findings on individual linguistic variation at the sentence level offer valuable guidance for developing more user-specific word embeddings and NLP systems.
@@ II-E2 Applications of Personalized Embeddings
Personalized word embeddings have proven beneficial across various NLP tasks. [charles_exploring_2020] demonstrated the potential of personalized embeddings in authorship attribution, showing that they can effectively capture unique writing styles. [hamid_deep_2015] introduced LSTM-RNN models for sentence embedding trained on user click-through data, underscoring the application of personalized embeddings in user-specific document retrieval systems that provide more relevant search results and tailored recommendations. Personalized embeddings can also leverage user history data: [doddapaneni_user_2024] proposed a User Embedding Module (UEM) that compresses user histories into embeddings to serve as prompts in language models. This approach yielded improvements in personalized language tasks, suggesting that personalized embeddings effectively capture user preferences and interests for customized recommendations and user profiling.
@@ III-A1 Simple Averaging and Pooling of Word Embeddings
  • Methods: Common methods include averaging the word embeddings, or using element-wise maximum (max-pooling), minimum (min-pooling), or sum operations. These approaches have been explored in various contexts, such as in a comparative study of word embedding models for hate speech detection [minni_comparison_2021].
  • Strengths: The main advantage of these methods is their simplicity and computational efficiency. They are easy to implement and require minimal computational resources, making them ideal for large-scale applications. For example, a study comparing bag-of-words approaches with sentence encoders found that simple averaging can be unexpectedly effective for certain tasks [c_evaluation_2018].
  • Limitations: However, these methods have notable limitations. Averaging or pooling word embeddings typically results in a loss of word order information, which limits their effectiveness in tasks where word order is significant. Additionally, they can be sensitive to outliers, as an uncommon word may disproportionately influence the resulting sentence embedding. To mitigate these issues, researchers have proposed more advanced techniques, such as using linear combinations of word embeddings to capture higher-level linguistic structures [lyndon_surprising_2019].
@@ III-A2 Recurrent Neural Network (RNN) based Approaches
  • Architectures: LSTM-RNNs and GRUs are commonly used for sentence embedding as they effectively capture long-range dependencies within sentences. These architectures surpass simple averaging methods in tasks that require deeper contextual understanding. For instance, a study on deep sentence embedding with LSTM networks demonstrated their capacity to detect keywords and allocate topics autonomously [hamid_deep_2015].
  • Training: RNN-based models can be trained using various approaches, including weakly supervised methods. One example involves training LSTM-RNNs with user click-through data from web search engines, which helps models learn effective representations for document retrieval [hamid_deep_2015].
  • Applications: RNN-generated sentence embeddings have been applied to numerous NLP tasks, such as document retrieval and semantic similarity measurement. For instance, using LSTM-RNN embeddings for document retrieval yielded notable performance improvements over traditional methods [hamid_deep_2015].
@@ III-A3 Transformer-based Sentence Encoders
Sentence-BERT (SBERT) has become a leading architecture for generating sentence embeddings, as highlighted by [a_knowledge_2020]. SBERT leverages pre-trained transformer models like BERT, fine-tuning them specifically for sentence-level tasks to generate efficient and effective sentence embeddings. Another notable model, SBERT-WK [bin_sbert-wk_2020], refines BERT-based word models to produce high-quality sentence representations. This model focuses on analyzing word representation evolution across BERT layers, applying geometric analysis to enhance sentence embeddings.
@@
Fine-tuning strategies are essential for adapting pre-trained language models to sentence embedding tasks. A common method is supervised fine-tuning on Natural Language Inference (NLI) and Semantic Textual Similarity (STS) datasets, as discussed in [fangxiaoyu_language-agnostic_2020]. This approach uses labeled sentence pairs with semantic relationship annotations (e.g., entailment, contradiction, similarity) to refine model performance. Knowledge distillation also proves effective; [nils_making_2020] demonstrates how a student model can be trained to mimic a teacher model, effectively transferring insights from a high-resource language to multiple target languages.
@@
Transformer-based sentence encoders achieve state-of-the-art results in various Semantic Textual Similarity (STS) tasks [j_bilingual_2019]. These models excel at capturing semantic relationships, making them highly effective for tasks requiring understanding of sentence meaning and similarity. Specifically, SBERT-WK demonstrates superior performance on STS benchmarks, further validating its robustness in sentence embedding applications [bin_sbert-wk_2020].
@@
Analyzing the evolution of word representations across transformer layers provides insights into how these models capture and process contextual information. Studies such as [r_lmfingerprints_2022] introduce methods to quantify and visualize contextualization across layers, often analyzing changes in word embeddings and examining the unique information captured by each layer. Additionally, [steven_analysing_2020] investigates word representations at both input and output stages, offering deeper insights into layer-wise contextual processing.
@@
The BeLLM model [xianming_bellm_2023] addresses the limitations of unidirectional attention in autoregressive large language models by introducing backward dependencies. By transforming specific attention layers to be bidirectional, BeLLM enhances semantic similarity tasks by considering both preceding and following context, thereby providing a more comprehensive contextual understanding.
@@ III-A4 Multilingual and Cross-lingual Sentence Embeddings
Knowledge distillation is an effective technique for aligning vector spaces across languages, facilitating cross-lingual understanding [nils_making_2020]. By training a student model to replicate the vector space of a monolingual teacher model, semantic knowledge from a high-resource language can be transferred to other languages, including those with limited resources.
@@
Dual encoder architectures and negative sampling are core techniques in translation ranking and bi-text mining tasks, as explored by [fangxiaoyu_language-agnostic_2020]. In these architectures, source and target sentences are encoded separately, with training focused on ranking true translation pairs higher than negative samples, thereby improving translation accuracy and relevance.
@@
The Language-agnostic BERT Sentence Embedding (LaBSE) model [fangxiaoyu_language-agnostic_2020] has demonstrated high performance in multilingual applications. LaBSE utilizes a combination of pre-training and dual-encoder fine-tuning to optimize translation ranking and produce language-agnostic sentence embeddings that excel across languages.
@@
Addressing language bias is essential to ensure fair and effective multilingual embeddings. [nils_making_2020] notes that some models may exhibit biases, favoring specific language combinations over others, which can impair performance in multilingual tasks. Techniques such as knowledge distillation are valuable for mitigating this bias and achieving more balanced cross-lingual representations.
@@
Handling low-resource languages remains challenging in multilingual NLP. [trevor_cross-lingual_2017] highlights that bilingual lexicons can provide an effective solution by enabling cross-lingual word embedding generation when text data is scarce. This approach enhances language model performance in low-resource environments, helping bridge the gap in multilingual understanding.
@@ III-B1 Methods for Combining Word Embeddings
A straightforward approach to creating document embeddings is to average the word embeddings of all words in the document. Although simple, this method can be highly effective and sometimes outperforms more complex models, particularly when word order is not critical [lyndon_surprising_2019]. Averaging captures semantic relationships by leveraging word co-occurrence patterns [lyndon_surprising_2019]. However, this method overlooks word order and may miss nuanced information in complex sentences [lyndon_surprising_2019].
Pooling methods, such as max-pooling or min-pooling, offer an alternative by selecting the most salient features from word embeddings. While these methods can highlight important features, they may still lose significant contextual information. More sophisticated techniques, like using recurrent neural networks (RNNs) with Long Short-Term Memory (LSTM) cells, process words sequentially, accumulating richer information as they traverse the document [hamid_deep_2015]. This method has shown effectiveness in tasks like document retrieval, as demonstrated in [hamid_deep_2015].
@@
Beyond simple averaging and pooling, compositional methods aim to capture interactions between words within a document. [mo_factor-based_2014] explores compositional embedding models that create representations for substructures within annotated sentences, addressing the challenge of representing a diverse range of sentences by building sentence embeddings from component word embeddings. Additionally, [lebret_word_2016] presents an approach that learns embeddings for phrases by combining word embeddings, providing richer phrase-level representations for documents.
@@ III-B2 Generative Topic Embedding Models
Integrating word embeddings into topic models, such as Latent Dirichlet Allocation (LDA), enhances the quality of learned topics. [shaohua_embedding_2016] proposes a generative topic embedding model that merges local word collocation patterns with global document-level topic patterns through variational inference. This approach yields topic embeddings and topic mixing proportions for each document, resulting in low-dimensional, continuous representations [shaohua_embedding_2016].
@@
Correlated topic models aim to capture relationships between topics. [guangxu_correlated_2017] introduces the Correlated Gaussian Topic Model (CGTM), which uses word embeddings to represent topics as multivariate Gaussian distributions. This structure captures semantic relatedness and correlations between words, improving topic coherence over traditional topic models [guangxu_correlated_2017].
@@
The Embedded Topic Model (ETM), described in [adji_topic_2019], combines word embeddings with traditional topic models to improve topic quality and predictive performance. Unlike LDA, ETM uses word and topic embeddings to model word probabilities within a topic, effectively handling large vocabularies and rare words. This model has demonstrated superior performance over LDA in settings with extensive vocabularies [adji_topic_2019].
@@ III-B3 Hierarchical Neural Language Models for Documents
Hierarchical neural models often use layered structures to capture different aspects of context. [nemanja_hierarchical_2015] presents a two-layer model: the upper layer models the temporal context of document sequences, assuming that temporally close documents are likely to be related, while the lower layer captures word sequence context within each document. This setup enables the model to learn joint representations for both documents and words within a shared feature space, which simplifies tasks like document tagging through nearest-neighbor search. This hierarchical structure can be extended with additional layers, such as a user layer, to incorporate user preferences for personalized recommendations.
@@
A key advantage of hierarchical neural models is their ability to incorporate temporal context. [nemanja_hierarchical_2015] shows how these models learn representations that account for temporal co-occurrences of documents within a stream. By optimizing the joint log-likelihood of document and word sequences, the model balances context at both levels, moving beyond traditional bag-of-words approaches. This temporal awareness supports a more dynamic understanding of language, resulting in improved performance in tasks like movie classification (MovieLens dataset) and click-through prediction (Yahoo News dataset).
@@
The joint representations of documents and words learned by hierarchical neural language models enable a range of applications. For document retrieval, the similarity between a query and document can be efficiently computed by comparing their vector representations in the shared embedding space. In recommendation systems, the model leverages user preferences, represented in an additional user layer, to generate personalized suggestions. For automatic tagging, the shared embedding space facilitates nearest-neighbor searches to assign relevant tags based on document vectors. [nemanja_hierarchical_2015] demonstrates the model’s superior performance on tasks like movie classification, underscoring its versatility in document retrieval and recommendation. The ability to capture both document- and word-level context makes these models highly applicable for text mining.
Sentence and Document Embeddings Cross-Lingual and Multi-Lingual Embeddings Cross-Lingual Word Embeddings [will_bilingual_2013, trevor_cross-lingual_2017] Challenges and Evaluation [takashi_unsupervised_2018, sebastian_survey_2017] Language-Agnostic Sentence Embeddings [nils_making_2020, fangxiaoyu_language-agnostic_2020] Document Embeddings Methods for Combining Word Embeddings [lyndon_surprising_2019, hamid_deep_2015, mo_factor-based_2014, lebret_word_2016] Generative Topic Embedding Models [shaohua_embedding_2016, guangxu_correlated_2017, adji_topic_2019] Hierarchical Neural Language Models [nemanja_hierarchical_2015] Sentence Embeddings Simple Averaging and Pooling [minni_comparison_2021, c_evaluation_2018, lyndon_surprising_2019] Recurrent Neural Networks [hamid_deep_2015] Transformer-based Sentence Encoders [a_knowledge_2020, bin_sbert-wk_2020, fangxiaoyu_language-agnostic_2020, nils_making_2020, r_lmfingerprints_2022, steven_analysing_2020, xianming_bellm_2023]
@@ III-C1 Cross-Lingual Word Embeddings: Methods and Applications
Cross-lingual word embeddings map words from different languages into a shared vector space, facilitating cross-lingual knowledge transfer. This approach is especially useful for low-resource languages, which often lack sufficient monolingual training data. Supervised methods typically use bilingual lexicons or word alignments from parallel corpora. For example, [will_bilingual_2013] introduces bilingual word embeddings learned from a large, unlabeled corpus, where machine translation word alignments are used to enforce translational equivalence. These embeddings, when integrated into a phrase-based machine translation system, show significant improvements in translation quality. Another supervised approach, explored in [trevor_cross-lingual_2017], leverages bilingual lexicons to enhance language models for low-resource languages, applying this technique to the language Yongning Na.
In contrast, unsupervised methods develop cross-lingual embeddings without explicit parallel data. [takashi_unsupervised_2018] proposes a multilingual neural language model that shares parameters across languages, jointly learning word embeddings in a shared space without parallel data or pre-training. This method proves effective even with limited monolingual data or domain differences across languages. For a comprehensive overview of contextual and cross-lingual embeddings, [qi_survey_2020] offers valuable insights.